# Existence and non-existence of consistent estimators in supercritical controlled branching processes

### v1

@@ 5 Proofs
Since $\varepsilon(z)=\Theta(z)$ and $\nu^{2}(z)=O(z)$, there exists a positive constant $C_{1}$ such that $\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})\leq C_{1}$, so that
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}U^{2}_{n}<\sum_{n=1}^{\infty}\frac{C_{1}}{n^{2}}<\infty.$
@@
Then by Theorem 5.3, $\frac{1}{n}\sum_{n=1}^{\infty}U_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\hat{m}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}m$.
@@
Let us now further assume that that there exist positive constants $c$ and $d$ such that $\sup_{z\geq 1}\left\{\frac{|\upsilon(z)|}{z}\right\}\leq c$ and $\sup_{z\geq 1}\Big{\{}\frac{\mathbbm{E}(\phi(z)-\varepsilon(z))^{4}}{z^{2}}\Big{\}}\leq d$, and that $\mathbbm{E}(\xi-m)^{4}$ is finite.
@@
(ii) Strong consistency of $\boldsymbol{\bar{\sigma}^{2}_{n}}$: In this case, we consider that $m$ is known. We introduce the estimator
@@
$V_{n}:=\tilde{\sigma}_{n}^{2}-\sigma^{2},$
for which we see that
$\mathbbm{E}(V_{n}|\mathcal{F}_{n-1})$ $=\varepsilon^{-1}(Z_{n-1})\cdot\Big{\{}\text{Var}(Z_{n}|\mathcal{F}_{n-1})-m^{2}\cdot\nu^{2}(Z_{n-1})\Big{\}}-\sigma^{2}$
$=\varepsilon^{-1}(Z_{n-1})\cdot\Big{\{}\mathbbm{E}\big{(}Z_{n}^{2}\big{|}\mathcal{F}_{n-1}\big{)}-m^{2}\cdot\varepsilon^{2}(Z_{n-1})-m^{2}\cdot\nu^{2}(Z_{n-1})\Big{\}}-\sigma^{2}$
$=\varepsilon^{-1}(Z_{n-1})\cdot\mathbbm{E}\Big{(}\big{(}Z_{n}-m\cdot\phi(Z_{n-1})\big{)}^{2}\Big{|}\mathcal{F}_{n-1}\Big{)}-\sigma^{2}$
$=\varepsilon^{-1}(Z_{n-1})\cdot\mathbbm{E}\Bigg{(}\sum_{i=1}^{\phi(Z_{n-1})}(\xi_{i}-m)^{2}+2\sum_{i=1}^{\phi(Z_{n-1})}\sum_{j=1}^{i-1}(\xi_{i}-m)(\xi_{j}-m)\Bigg{|}\mathcal{F}_{n-1}\Bigg{)}-\sigma^{2}$
$=\varepsilon^{-1}(Z_{n-1})\cdot\sigma^{2}\cdot\mathbbm{E}\Big{(}\phi(Z_{n-1})\Big{|}\mathcal{F}_{n-1}\Big{)}-\sigma^{2}$
$=0,$
and, since $\mathbbm{E}V_{n}=\mathbbm{E}\big{(}\mathbbm{E}(V_{n}|\mathcal{F}_{n-1})\big{)}=0$, and given that $\varepsilon(z)=\Theta(z)$ and $\nu^{2}(z)=O(z)$, we have
$\mathbbm{E}|V_{n}|$ $\leq\mathbbm{E}\Bigg{|}\frac{(Z_{n}-m\cdot\varepsilon(Z_{n-1}))^{2}}{\varepsilon(Z_{n-1})}\Bigg{|}+\mathbbm{E}\Bigg{|}\frac{m^{2}\cdot\nu^{2}(Z_{n-1})}{\varepsilon(Z_{n-1})}\Bigg{|}+\sigma^{2}$
$=\mathbbm{E}V_{n}+2m^{2}\cdot\mathbbm{E}\Bigg{(}\frac{\nu^{2}(Z_{n-1})}{\varepsilon(Z_{n-1})}\Bigg{)}+2\sigma^{2}$
<<$\infty$.< ∞ .
@@
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}V_{n}^{2}=\sum_{n=1}^{\infty}\frac{1}{n^{2}}\Big{(}\mathbbm{E}\tilde{\sigma}_{n}^{4}-\sigma^{4}\Big{)}<\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}\tilde{\sigma}_{n}^{4}.$
We can see that
$\mathbbm{E}\big{(}\tilde{\sigma}_{n}^{4}\big{|}\mathcal{F}_{n-1}\big{)}$ $=\varepsilon^{-2}(Z_{n-1})\cdot\bigg{(}\mathbbm{E}\Big{[}\big{(}Z_{n}-m\cdot\varepsilon(Z_{n-1})\big{)}^{4}\Big{|}\mathcal{F}_{n-1}\Big{]}+m^{4}\nu^{4}(Z_{n-1})-2m^{2}\nu^{2}(Z_{n-1})\cdot\text{Var}(Z_{n}|\mathcal{F}_{n-1})\bigg{)}$
$=\varepsilon^{-2}(Z_{n-1})\cdot\mathbbm{E}\Big{[}\big{(}Z_{n}-m\cdot\varepsilon(Z_{n-1})\big{)}^{4}\Big{|}\mathcal{F}_{n-1}\Big{]}-2m^{2}\sigma^{2}\cdot\frac{\nu^{2}(Z_{n-1})}{\varepsilon(Z_{n-1})}-m^{4}\cdot\frac{\nu^{4}(Z_{n-1})}{\varepsilon^{2}(Z_{n-1})}$
@@
where, by expanding and simplifying, we can find
$\mathbbm{E}\Big{[}\big{(}Z_{n}-m\cdot\varepsilon(Z_{n-1})\big{)}^{4}\Big{|}\mathcal{F}_{n-1}\Big{]}$ $=m^{4}\mathbbm{E}(\phi(Z_{n-1})-\varepsilon(Z_{n-1}))^{4}+6\sigma^{2}m^{2}\upsilon(Z_{n-1})-12m^{2}\sigma^{2}\varepsilon(Z_{n-1})\nu^{2}(Z_{n-1})$
@@
$\mathbbm{E}\big{(}\tilde{\sigma}_{n}^{4}\big{|}\mathcal{F}_{n-1}\big{)}$ $<\frac{m^{4}\mathbbm{E}(\phi(Z_{n-1})-\varepsilon(Z_{n-1}))^{4}}{\varepsilon^{2}(Z_{n-1})}+\frac{6\sigma^{2}m^{2}\upsilon(Z_{n-1})}{\varepsilon^{2}(Z_{n-1})}$
@@
Given (8) and our assumptions on the bounds on $\mathbbm{E}(\phi(Z_{n-1})-\varepsilon(Z_{n-1}))^{4}$ and $\mathbbm{E}(\xi-m)^{4}$, there will therefore exist a positive constant $C_{2}$ such that $\mathbbm{E}\big{(}\tilde{\sigma}_{n}^{4}\big{|}\mathcal{F}_{n-1}\big{)}\leq C_{2}$. Hence
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}V_{n}^{2}<\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}\big{(}\mathbbm{E}\big{(}\tilde{\sigma}_{n}^{4}\big{|}\mathcal{F}_{n-1}\big{)}\big{)}<\sum_{n=1}^{\infty}\frac{C_{2}}{n^{2}}<\infty.$
Then by Theorem 5.3, $\frac{1}{n}\sum_{n=1}^{\infty}V_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\bar{\sigma}^{2}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma^{2}$.
@@
$\hat{\sigma}^{2}_{n}=\bar{\sigma}^{2}_{n}+\underbrace{\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}}_{\text{I}}+\underbrace{\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\varepsilon(Z_{k-1})}_{\text{II}}+\underbrace{\frac{m^{2}-\tilde{m}^{2}_{n}}{n}\sum_{k=1}^{n}\frac{\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})},}_{\text{III}}$ (19)
@@
    It follows from Chebyshev’s inequality that
    $\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$
    if we can find that
    $\mathbbm{E}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0$ (20)
    and
    $\text{Var}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0.$ (21)
    We can find
    $\mathbbm{E}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{|}\mathcal{F}_{n-1}\bigg{)}$
    $=\frac{2}{n}\sum_{k=1}^{n-1}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\cdot\mathbbm{E}(m-\tilde{m}_{n}|\mathcal{F}_{n-1})+\frac{2}{n}\cdot\mathbbm{E}\big{(}(m-\tilde{m}_{n})(Z_{n}-m\cdot\varepsilon(Z_{n-1}))\big{|}\mathcal{F}_{n-1}\big{)}$
    $=-\frac{2}{n\cdot\varepsilon(Z_{n-1})}\cdot\text{Var}\big{(}Z_{n}\big{|}\mathcal{F}_{n-1}\big{)}$
    $=-\frac{2\sigma^{2}}{n}-\frac{2m^{2}\cdot\nu^{2}(Z_{n-1})}{n\cdot\varepsilon(Z_{n-1})}.$
    Since $\varepsilon(z)=\Theta(z)$ and $\nu^{2}(z)=O(z)$, there exists a positive constant $C_{3}$ such that $\frac{\nu^{2}(z)}{\varepsilon(z)}\leq C_{3}$ for all $z$, so that
    $-\frac{2\sigma^{2}}{n}-\frac{2m^{2}C_{3}}{n}\leq\mathbbm{E}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\leq-\frac{2\sigma^{2}}{n}.$
    Since $\mathbbm{E}\Big{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\Big{|}\mathcal{F}_{n-1}\Big{)}$ is squeezed between two functions that both converge to zero as $n\to\infty$, (20) follows as a consequence, and it will also be the case that
    $\text{Var}\bigg{(}\mathbbm{E}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0.$
    Consequently, since
    $\text{Var}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{)}$ $=\mathbbm{E}\bigg{(}\text{Var}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\bigg{)}$
    $\quad+\text{Var}\bigg{(}\mathbbm{E}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\bigg{)},$
    (21) will follow if we show that
    $\mathbbm{E}\bigg{(}\text{Var}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0.$ (22)
    Indeed,
@@
    $\text{Var}\bigg{(}\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{|}\mathcal{F}_{n-1}\bigg{)}$
@@
    $=\frac{4}{n^{2}}\bigg{(}\sum_{k=1}^{n-1}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{)}^{2}\cdot\text{Var}\big{(}m-\tilde{m}_{n}\big{|}\mathcal{F}_{n-1}\big{)}+\frac{4}{n^{2}}\cdot\text{Var}\big{(}(m-\tilde{m}_{n})(Z_{n}-m\cdot\varepsilon(Z_{n-1}))\big{|}\mathcal{F}_{n-1}\big{)}$
    $=\frac{4}{n^{2}}\bigg{(}\sum_{k=1}^{n-1}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{)}^{2}\cdot\text{Var}\big{(}\tilde{m}_{n}\big{|}\mathcal{F}_{n-1}\big{)}+\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\cdot\text{Var}\big{(}(Z_{n}-m\cdot\varepsilon(Z_{n-1}))^{2}\big{|}\mathcal{F}_{n-1}\big{)}$
    $=\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\bigg{(}\sum_{k=1}^{n-1}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{)}^{2}\cdot\text{Var}(Z_{n}|\mathcal{F}_{n-1})+\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\cdot\mathbbm{E}\big{(}\big{(}Z_{n}-m\cdot\varepsilon(Z_{n-1})\big{)}^{4}\big{|}\mathcal{F}_{n-1}\big{)}$
    $\quad+\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\cdot\big{(}\text{Var}(Z_{n}|\mathcal{F}_{n-1})\big{)}^{2},$
    so, since
    $\mathbbm{E}\bigg{(}\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\bigg{(}\sum_{k=1}^{n-1}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}\bigg{)}^{2}\cdot\text{Var}(Z_{n}|\mathcal{F}_{n-1})\bigg{)}$
    $=\mathbbm{E}\bigg{(}\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\mathbbm{E}\bigg{(}\sum_{k=1}^{n-1}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}^{2}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\cdot\text{Var}(Z_{n}|\mathcal{F}_{n-1})\bigg{)}$
    $=\mathbbm{E}\bigg{(}\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\sum_{k=1}^{n-1}\text{Var}(Z_{k}|\mathcal{F}_{k-1})\cdot\text{Var}(Z_{n}|\mathcal{F}_{n-1})\bigg{)}$
    $\leq\frac{4(a\sigma^{2}+bm^{2})^{2}}{n^{2}}\cdot\sum_{k=1}^{n-1}\mathbbm{E}\bigg{(}\frac{Z_{k-1}\cdot Z_{n-1}}{\varepsilon^{2}(Z_{n-1})}\bigg{)}$
    $\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
    we can see from (5.2) that
    $\mathbbm{E}\bigg{(}\frac{4}{n^{2}\cdot\varepsilon^{2}(Z_{n-1})}\cdot\mathbbm{E}\big{(}\big{(}Z_{n}-m\cdot\varepsilon(Z_{n-1})\big{)}^{4}\big{|}\mathcal{F}_{n-1}\big{)}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
    and, since $\varepsilon(z)=\Theta(z)$ and $\nu^{2}(z)=O(z)$,

### v2

@@ 6 Proofs of consistency results
    Since $\varepsilon(z)>0$ for all $z$ by assumption, and since $\lim_{z\to\infty}\varepsilon(z)=\infty$ by the supercriticality of the process, we can reason that $\{\varepsilon(z)\}_{z\in\mathbbm{N}_{0}}$ has its minimum bounded away from zero. Hence there exists a positive constant $C_{1}$ such that $\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})\leq C_{1}$, so that
    $\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})<\sum_{n=1}^{\infty}\frac{C_{1}}{n^{2}}<\infty.$
    Then by Corollary 6.2, $\frac{1}{n}\sum_{k=1}^{n}U_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\hat{m}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}m$.
  2. (b)
    In this case, there are two main differences with (a): (i) due to the existence of some $z$ such that $\varepsilon(z)=0$, the index set $I^{\varepsilon}_{n}$ is no longer equivalent to $\{1,\dots,n\}$, and (ii) if $\varepsilon(0)=0$, then 0 is an absorbing state, which implies $\mathbbm{P}(Z_{n}\to\infty)<1$, and we have to carefully deal with the restriction to the set of unbounded growth.
    We deal with both (i) and (ii) by defining a modified process that satisfies case (a) and has the same asymptotic properties as the original process on $\{Z_{n}\to\infty\}$. Without loss of generality, we assume that $\varepsilon(z_{0})>0$ (otherwise, a similar argument can be applied).
    We let $L_{\varepsilon}=\{z\in\mathbbm{N}_{0}:\varepsilon(z)=0\}$ and $\{Z^{\uparrow}_{n},z_{0}\}$ be a process with
    $\xi^{\uparrow}\stackrel{{\scriptstyle d}}{{=}}\xi\quad\text{and}\quad\phi^{\uparrow}(z)\stackrel{{\scriptstyle d}}{{=}}\begin{cases}z_{0}\;\;\text{a.s.,}&\text{if}\;\;z\in L_{\varepsilon}\\ \phi(z),&\text{otherwise.}\end{cases}$
    This implies that, for all $z$ such that $\varepsilon(z)>0$,
    $(Z^{\uparrow}_{n}|Z^{\uparrow}_{n-1}=z)\stackrel{{\scriptstyle d}}{{=}}(Z_{n}|Z_{n-1}=z),$
    but $\{Z^{\uparrow}_{n},z_{0}\}$ has $\varepsilon^{\uparrow}(z)>0$ for all $z\in\mathbbm{N}_{0}$, and therefore constitutes a process satisfying case (a). Hence
@@
    $\frac{1}{n}\sum_{k=1}^{n}U^{\uparrow}_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0\;\;\text{as}\;\;n\to\infty,\quad\text{for}\;\;U^{\uparrow}_{k}:=\frac{Z^{\uparrow}_{k}}{\varepsilon^{\uparrow}(Z^{\uparrow}_{k-1})}-m.$
    We define $T^{\uparrow}_{0}$, respectively $T_{0}$, as the last time $\{Z^{\uparrow}_{n},z_{0}\}$, resp. $\{Z_{n},z_{0}\}$, visits a state in $L_{\varepsilon}$ before escaping to infinity, with the convention that $T^{\uparrow}_{0}=0$, resp. $T_{0}=0$, if the process never visits $L_{\varepsilon}$. Because $\mathbbm{P}(Z^{\uparrow}_{n}\to\infty)=1$, $T^{\uparrow}_{0}$ is almost surely finite. Similarly, on $\{Z_{n}\to\infty\}$, $T_{0}$ is almost surely finite. We then have
    $\{Z^{\uparrow}_{n},z_{0}\}_{n\geq T^{\uparrow}_{0}}\stackrel{{\scriptstyle d}}{{=}}\{Z_{n},z_{0}\}_{n\geq T_{0}}\quad\text{on}\quad\{Z_{n}\to\infty\},$
    so that
    $\lim_{n\to\infty}\frac{1}{n}\sum_{k=T^{\uparrow}_{0}}^{n}U^{\uparrow}_{k}=\lim_{n\to\infty}\frac{1}{n}\sum_{k=T_{0}}^{n}U_{k}\quad\text{a.s.\ on}\;\{Z_{n}\to\infty\}.$ (28)
    Since there are an a.s. finite number of generations (i.e., terms in the sum) before $T^{\uparrow}_{0}$ and $T_{0}$, we have
    $\frac{1}{n}\sum_{k=1}^{n}U^{\uparrow}_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0\implies\frac{1}{n}\sum_{k=T^{\uparrow}_{0}}^{n}U^{\uparrow}_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0\quad\text{and}\quad\frac{1}{n}\sum_{k=T_{0}}^{n}U_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0\implies\frac{1}{|I^{\varepsilon}_{n}|}\sum_{k\in I^{\varepsilon}_{n}}U_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0.$ (29)
    By combining (28) and (29), it follows that $\hat{m}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}m$ on the event $\{Z_{n}\to\infty\}$.
@@
With the consistency of $\hat{m}_{n}$ shown, it remains to prove the consistency of $\bar{\sigma}^{2}_{n}$ and $\hat{\sigma}^{2}_{n}$. Accordingly, we now further assume that that there exist positive constants $c$ and $d$ such that $\sup_{z\geq 1}\left\{\frac{|\iota(z)|}{z}\right\}\leq c$ and $\sup_{z\geq 1}\Big{\{}\frac{\mathbbm{E}(\phi(z)-\varepsilon(z))^{4}}{z^{2}}\Big{\}}\leq d$, and that $\mathbbm{E}(\xi-m)^{4}$ is finite. In addition, to avoid repeating ourselves, we hereafter also assume that $\varepsilon(z)>0$ for all $z\in\mathbbm{N}_{0}$, so that our two estimators simplify to
$\bar{\sigma}^{2}_{n}:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-m\cdot\varepsilon(Z_{k-1}))^{2}-m^{2}\cdot\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}$
and
$\hat{\sigma}^{2}_{n}:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-\tilde{m}_{n}\cdot\varepsilon(Z_{k-1}))^{2}-\tilde{m}_{n}^{2}\cdot\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}$
We omit the extension to the case where $\varepsilon(z)=0$ for some $z\in\mathbbm{N}_{0}$, which follows the same argument as in in case (b) above.
@@
(ii) Strong consistency of $\boldsymbol{\bar{\sigma}^{2}_{n}}$: In this case, we consider $m$ to be known. We introduce the estimator
@@
$U_{n}:=\tilde{\sigma}_{n}^{2}-\sigma^{2},$
for which we use (6) to see that
$\mathbbm{E}(U_{n}|\mathcal{F}_{n-1})=\varepsilon^{-1}(Z_{n-1})\cdot\Big{\{}\text{Var}(Z_{n}|\mathcal{F}_{n-1})-m^{2}\cdot\nu^{2}(Z_{n-1})\Big{\}}-\sigma^{2}=0.$
Thus $\mathbbm{E}U_{n}=\mathbbm{E}\big{(}\mathbbm{E}(U_{n}|\mathcal{F}_{n-1})\big{)}=0$, and given that $\nu^{2}(Z_{n-1})/\varepsilon(Z_{n-1})<C$ by (27), we have
$\mathbbm{E}|U_{n}|$ $\leq\mathbbm{E}\Bigg{|}\frac{(Z_{n}-m\cdot\varepsilon(Z_{n-1}))^{2}}{\varepsilon(Z_{n-1})}\Bigg{|}+\mathbbm{E}\Bigg{|}\frac{m^{2}\cdot\nu^{2}(Z_{n-1})}{\varepsilon(Z_{n-1})}\Bigg{|}+\sigma^{2}$
$=\mathbbm{E}U_{n}+2m^{2}\cdot\mathbbm{E}\Bigg{(}\frac{\nu^{2}(Z_{n-1})}{\varepsilon(Z_{n-1})}\Bigg{)}+2\sigma^{2}$
$<\infty.$
@@
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}U_{n}^{2}=\sum_{n=1}^{\infty}\frac{1}{n^{2}}\Big{(}\mathbbm{E}\tilde{\sigma}_{n}^{4}-\sigma^{4}\Big{)}<\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}\tilde{\sigma}_{n}^{4}.$
We have that
$\mathbbm{E}\big{(}\tilde{\sigma}_{n}^{4}\big{|}\mathcal{F}_{n-1}\big{)}$ $=\varepsilon^{-2}(Z_{n-1})\cdot\mathbbm{E}\Big{[}\big{(}Z_{n}-m\cdot\varepsilon(Z_{n-1})\big{)}^{4}\Big{|}\mathcal{F}_{n-1}\Big{]}-2m^{2}\sigma^{2}\cdot\frac{\nu^{2}(Z_{n-1})}{\varepsilon(Z_{n-1})}-m^{4}\cdot\frac{\nu^{4}(Z_{n-1})}{\varepsilon^{2}(Z_{n-1})}$
@@
and by a lengthy yet elementary expansion can find that
$\mathbbm{E}\Big{[}\big{(}Z_{n}-m\cdot\varepsilon(Z_{n-1})\big{)}^{4}\Big{|}\mathcal{F}_{n-1}\Big{]}$ $=m^{4}\mathbbm{E}(\phi(Z_{n-1})-\varepsilon(Z_{n-1}))^{4}+6\sigma^{2}m^{2}\iota(Z_{n-1})$
$\quad-12m^{2}\sigma^{2}\varepsilon(Z_{n-1})\nu^{2}(Z_{n-1})$
@@
$\mathbbm{E}\big{(}\tilde{\sigma}_{n}^{4}\big{|}\mathcal{F}_{n-1}\big{)}$ $<\frac{m^{4}\mathbbm{E}(\phi(Z_{n-1})-\varepsilon(Z_{n-1}))^{4}}{\varepsilon^{2}(Z_{n-1})}+\frac{6\sigma^{2}m^{2}\iota(Z_{n-1})}{\varepsilon^{2}(Z_{n-1})}$
@@
Given (6) and our assumptions on the bounds on $|\iota(Z_{n-1})|$, $\mathbbm{E}(\phi(Z_{n-1})-\varepsilon(Z_{n-1}))^{4}$ and $\mathbbm{E}(\xi-m)^{4}$, there will therefore exist a positive constant $C_{1}$ such that $\mathbbm{E}\big{(}\tilde{\sigma}_{n}^{4}\big{|}\mathcal{F}_{n-1}\big{)}\leq C_{1}$. Hence
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}U_{n}^{2}<\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}\big{(}\mathbbm{E}\big{(}\tilde{\sigma}_{n}^{4}\big{|}\mathcal{F}_{n-1}\big{)}\big{)}<\sum_{n=1}^{\infty}\frac{C_{1}}{n^{2}}<\infty.$
Consequently, by Corollary 6.2, $\frac{1}{n}\sum_{k=1}^{n}U_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\bar{\sigma}^{2}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma^{2}$.
@@
$\hat{\sigma}^{2}_{n}=\bar{\sigma}^{2}_{n}+\underbrace{\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}}_{\text{I}}+\underbrace{\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\varepsilon(Z_{k-1})}_{\text{II}}+\underbrace{\frac{m^{2}-\tilde{m}^{2}_{n}}{n}\sum_{k=1}^{n}\frac{\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})},}_{\text{III}}$
@@
    Recalling that $\tilde{m}_{n}:=Z_{n}/\varepsilon(Z_{n-1})$, we can decompose
@@
    $\hskip-17.07182pt\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}\big{(}Z_{k}-m\cdot\varepsilon(Z_{k-1})\big{)}=-2\underbrace{\sqrt{\frac{\sum_{k=1}^{n}\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}}}_{\text{(a)}}\cdot\underbrace{\frac{Z_{n}-m\varepsilon(Z_{n-1})}{n^{1/4}\cdot\sqrt{\varepsilon(Z_{n-1})}}}_{\text{(b)}}\cdot\underbrace{\frac{\sum_{k=1}^{n}(Z_{k}-m\varepsilon(Z_{k-1}))}{n^{3/4}\sqrt{\sum_{k=1}^{n}\varepsilon(Z_{k-1})}}}_{\text{(c)}},$
    and consider each of the components $(a)$, $(b)$ and $(c)$ in turn.
    1. (a)
      Given (6), $\varepsilon(z)\leq az$. Additionally, for $t$ such that $1<t<\liminf_{z\to\infty}\tau(z)$, there exists $N>0$ such that, for $z\geq N$, $\varepsilon(z)>\frac{t}{m}z$. Therefore, on $\{Z_{n}\to\infty\}$,
      $\lim_{n\to\infty}\frac{\sum_{k=1}^{n}\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}\leq\lim_{n\to\infty}\frac{am}{t}\cdot\frac{\sum_{k=1}^{n}Z_{k-1}}{Z_{n-1}}\quad\text{a.s.}$ (31)
      Then, as a result of Lemma 6.4, for any $s$ such that $1<s<t$, we have that
      $\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sqrt{\frac{\sum_{k=1}^{n}\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}}\leq\sqrt{\frac{t}{am(1-s^{-1})}}\Bigg{)}=1,$
      that is, $\sqrt{\frac{\sum_{k=1}^{n}\varepsilon(Z_{k-1})}{\varepsilon(Z_{n-1})}}$ is asymptotically bounded by a constant in probability.
    2. (b)
      Given that
      $\mathbbm{E}\bigg{(}\frac{Z_{n}-m\varepsilon(Z_{n-1})}{n^{1/4}\cdot\sqrt{\varepsilon(Z_{n-1})}}\bigg{|}\mathcal{F}_{n-1}\bigg{)}=0,$
      and given that, by (6) and (27),
@@
      $\text{Var}\bigg{(}\frac{Z_{n}-m\varepsilon(Z_{n-1})}{n^{1/4}\cdot\sqrt{\varepsilon(Z_{n-1})}}\bigg{|}\mathcal{F}_{n-1}\bigg{)}=\frac{\sigma^{2}\cdot\varepsilon(Z_{n-1})+m^{2}\cdot\nu^{2}(Z_{n-1})}{\sqrt{n}\cdot\varepsilon(Z_{n-1})}\leq\frac{\sigma^{2}+m^{2}C}{\sqrt{n}}=O\bigg{(}\frac{1}{\sqrt{n}}\bigg{)},$
      we see that